OpenStudy (anonymous):

4. Gavin and Seiji both worked hard over the summer. Together, they earned a total of $425. Gavin earned $25 more than Seiji. How much did each of them earn? (a) Write a system of two equations with two variables to model this problem. (b) Use substitution or the elimination method to solve the system. (c) Graph both equations. (d) Answer the question. I Just Need Help With (c) Medal & Fan!

OpenStudy (anonymous):

i dont know how to graph this

OpenStudy (boldjon):

s

OpenStudy (anonymous):

?

OpenStudy (boldjon):

well here u go S+g=425 g=s+25 subsisute s+s+25=425 2s+25=425 subtract 25 2s=400 divide 2 s=200 subsisute g=200+25 g=225 seiji=200 gavin=225

OpenStudy (anonymous):

thats off a website

OpenStudy (boldjon):

and that's the answer

OpenStudy (anonymous):

i know what website it is i need to solve for c

OpenStudy (anonymous):

i have already solved for everything else i just need (c)

OpenStudy (boldjon):

does this help Let the money Seiji got be x and Gavin, y. Total is 425. So, x + y = 425 Gavin got 25 more than Seiji. So, y = x + 25 I will use substitution method. Substitute y = x + 25 in x + y = 425. x + x + 25 = 425 2x + 25 = 425 2x = 400 x=$200 So, x is Seiji and he got $200. For Gavin, substitute x in the equation of y. y = x + 25 y = 200 + 25 y = $225 So, y is Gavin and he got $225.

OpenStudy (anonymous):

no